# Topological Sequence Analysis of Genomes: Category Approaches

### v1

@@ 2.1 Category theory
    $h\circ(g\circ f)=(h\circ g)\circ f;$
@@
    $\mathrm{id}_{Y}\circ f=f=f\circ\mathrm{id}_{X}.$
@@ 2.2 Persistent Homology
$H_{k}^{a,b}:=\mathrm{Im}\left(H_{k}(X_{a})\to H_{k}(X_{b})\right),$
@@ 3.1 The resolution category
$\sigma=(x_{j+k},x_{j+k+1},\dots,x_{j+k+|\sigma|-1})\quad\text{and}\quad i=j+k.$
@@
$\mathrm{ob}(\mathcal{C}_{\xi})=\coprod_{k\geq 0}\mathcal{S}_{\xi}(k).$
@@
The objects of $\mathcal{C}_{\xi}$ are pairs $(\sigma,i)$, where $\sigma$ is a contiguous subword of $\xi$, and $i$ is the starting position of this subword in $\xi$ (positions start from 00). For example,
@@
  • $(\texttt{abc},0)$ denotes abc at position 00.
@@
  • $(\texttt{a},0)\to(\texttt{abc},0)$, since a is a prefix of abc starting at 00.
@@ 3.2 Distances on resolution category
The Levenshtein distance is a widely used distance that quantifies the minimum number of edit operations (insertions, deletions, and substitutions) required to transform one sequence into another. The LCS distance, as the name suggests, is based on the length of the longest common subsequence shared by two sequences, providing a measure of their structural similarity. The Hamming distance, particularly suitable for sequences of equal length, counts the number of positions at which the corresponding elements differ. However, these classical distances primarily focus on the intrinsic differences between two sequences themselves. They are not designed to capture the differences in the occurrences or contextual roles of two sequences within a larger sequence.
@@
$(\sigma_{1},i_{1})\leftarrow(\sigma_{0},i_{0})\rightarrow(\sigma_{2},i_{2})$
@@
$d_{\mathrm{int}}((\sigma_{1},i_{1}),(\sigma_{2},i_{2})):=1-\frac{L}{\min\{\mathrm{length}(\sigma_{1}),\mathrm{length}(\sigma_{2})\}}.$
This distance lies in $[0,1]$, and takes the value 00 if and only if the shorter of the two subsequences is fully contained in a common intersection object. It increases as the shared structure between the two objects becomes smaller.
@@
$d_{\mathrm{pos}}((\sigma_{1},i_{1}),(\sigma_{2},i_{2}))=|i_{1}-i_{2}|.$
@@
$S(\sigma)=\{\,(\sigma,i)\in\mathrm{Ob}(\mathcal{C}_{\xi})\mid\text{$\sigma$ occurs in $\xi$ starting at position $i$}\,\}.$
@@
$d_{\mathrm{freq}}(\sigma_{1},\sigma_{2})=\left|f_{\xi}(\sigma_{1})-f_{\xi}(\sigma_{2})\right|.$
@@
$d((\sigma_{1},i_{1}),(\sigma_{2},i_{2}))=\alpha\cdot d_{\text{LCS}}(\sigma_{1},\sigma_{2})+\beta\cdot d_{\text{pos}}((\sigma_{1},i_{1}),(\sigma_{2},i_{2})),$
@@ 3.3 Substructure complexes in the resolution category
Assume that a distance function $d$ is defined on the set of objects of $\mathcal{C}_{\xi}$. Let $U\subseteq\operatorname{Ob}(\mathcal{C}_{\xi})$ be a finite collection of objects. For each real number $r\in\mathbb{R}$, we define a $\Delta$-complex $\mathcal{K}_{r}(U)$ as follows: for $p\geq 1$, a $p$-simplex in $\mathcal{K}_{r}(U)$ is a subset $\{\sigma_{0},\dots,\sigma_{p}\}\subseteq U$ of $p+1$ objects such that $d(\sigma_{i},\sigma_{j})\leq r$ for all $0\leq i,j\leq p$. And the vertex set of $\mathcal{K}_{r}(U)$ is $U/\sim$. Here, $\sim$ represents equivalence of objects at the same position, i.e., $(\sigma_{1},i)\sim(\sigma_{2},i)$ for any $\sigma_{1}$ and $\sigma_{2}$. This complex is called the substructure complex associated with $U$.
Consider the case where $U=S(\sigma)$. In this setting, $\mathcal{K}_{r}(U)$ coincides with the Vietoris–Rips complex of the metric subspace $(U,d)$ at scale parameter $r$. As $r$ increases, this construction defines a filtration $\{\mathcal{K}_{r}(U)\}_{r\in\mathbb{R}}$, which gives rise to persistent homology. The resulting persistence module captures the multi-scale topological features of the object collection $U$ within $\mathcal{C}_{\xi}$.
@@
$H_{n}^{a,b}(U)\coloneqq{\mathrm{im}\hskip 1.00006pt}\left(f^{a,b}_{n}:H_{n}(\mathcal{K}_{a}(U))\to H_{n}(\mathcal{K}_{b}(U))\right),\quad n\geq 0.$
@@
$d_{\mathrm{pos}}((\sigma_{1},i_{1}),(\sigma_{2},i_{2}))=|i_{1}-i_{2}|.$
@@
$\begin{pmatrix}&(C,1)&(G,3)&(C,5)&(G,8)\\ (C,1)&0&2&4&7\\ (G,3)&2&0&2&5\\ (C,5)&4&2&0&3\\ (G,8)&7&5&3&0\\ \end{pmatrix}$
@@
$H_{n}(\mathcal{K}_{2}(U))=\begin{cases}\mathbb{K}\oplus\mathbb{K},&\text{if }n=0,\\ 0,&\text{otherwise}.\end{cases}$
@@
$\mathcal{K}_{0}(U)$ =$=\{\{v_{1}\},\{v_{2}\},\{v_{3}\},\{v_{4}\}\},$
$\mathcal{K}_{1}(U)$ =$=\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{1},v_{3}\},\{v_{3},v_{4}\}\},$
$\mathcal{K}_{2}(U)$ =$=\{\{v_{1},v_{2},v_{3}\}\}.$
@@
$H_{n}^{2,4}(U)=\begin{cases}\mathbb{K},&\text{if }n=0,\\ 0,&\text{otherwise}.\end{cases}$
@@
Let $U=\{A,AG\}$ be the subset of nucleic acid patterns of interest. The positions of $A$ and $AG$ in $\xi$ are as follows: $A$ appears at positions 00, $6$, and $7$, and $AG$ appears at position $7$. We associate these occurrences with labeled objects as follows: $(A,0)$, $(A,6)$, $(A,7)$, and $(AG,7)$. For simplicity, we denote these objects as $w_{1}=(A,0)$, $w_{2}=(A,6)$, $w_{3}=(A,7)$, and $w_{4}=(AG,7)$.
@@
$\begin{pmatrix}&(A,0)&(A,6)&(A,7)&(AG,7)\\ (A,0)&0&6&7&7\\ (A,6)&6&0&1&1\\ (A,7)&7&1&0&0\\ (AG,7)&7&1&0&0\\ \end{pmatrix}$
@@
$\mathcal{K}_{0}(U)$ =$=\{\{w_{1}\},\{w_{2}\},\{w_{3}\}=\{w_{4}\}\},$
$\mathcal{K}_{1}(U)$ =$=\{\{w_{2},w_{3}\},\{w_{2},w_{4}\},\{w_{3},w_{4}\}\},$
$\mathcal{K}_{2}(U)$ =$=\{\{w_{2},w_{3},w_{4}\}\}.$
@@
$H_{n}(\mathcal{K}_{1}(U))=\begin{cases}\mathbb{K}\oplus\mathbb{K},&\text{if }n=0,\\ \mathbb{K},&\text{if }n=1,\\ 0,&\text{otherwise}.\end{cases}$
@@
$\mathcal{K}_{0}(U)$ =$=\{\{w_{1}\},\{w_{2}\},\{w_{3}\}=\{w_{4}\}\},$
$\mathcal{K}_{1}(U)$ =$=\{\{w_{1},w_{2}\},\{w_{2},w_{3}\},\{w_{2},w_{4}\},\{w_{3},w_{4}\}\},$
$\mathcal{K}_{2}(U)$ =$=\{\{w_{2},w_{3},w_{4}\}\}.$
@@
$H_{n}^{1,6}(U)=\begin{cases}\mathbb{K},&\text{if }n=0,1,\\ 0,&\text{otherwise}.\end{cases}$
@@ 4.1 CTSA for predictions of protein-nucleic acid binding affinities
For nucleic acids, we validate CTSA in DNA sequences to assess its capability to reveal essential interaction patterns relevant to protein binding. CTSA provides a topological realization of DNA by introducing resolution categories that reflect the intricate connections of the sequence within its k-mer spaces. Persistent homology is then applied across these categories to extract topological features, namely, persistent Betti numbers, which are essential for CTSA embeddings to capture both local and global structure of sequence.
@@ 5 Concluding remarks
6 Acknowledgments
This work was supported in part by NIH grants R01GM126189, R01AI164266, and R35GM148196, National Science Foundation grants DMS2052983, DMS-1761320, and IIS-1900473, NASA grant 80NSSC21M0023, Michigan State University Research Foundation, and Bristol-Myers Squibb 65109. Jian was also supported by the Natural Science Foundation of China (NSFC Grant No. 12401080) and the start-up research fund from Chongqing University of Technology.

### v2

# Topological sequence analysis of genomes: category theory approaches

@@ 2.1 Category theory
    $h\circ(g\circ f)=(h\circ g)\circ f;$ (1)
@@
    $\mathrm{id}_{Y}\circ f=f=f\circ\mathrm{id}_{X}.$ (2)
@@ 2.2 Persistent Homology
$H_{k}^{a,b}:=\mathrm{Im}\left(H_{k}(X_{a})\to H_{k}(X_{b})\right),$ (4)
@@ 3.1 The resolution category
$\sigma=(x_{j+k},x_{j+k+1},\dots,x_{j+k+|\sigma|-1})\quad\text{and}\quad i=j+k.$ (7)
@@
$\mathrm{ob}(\mathcal{C}_{\xi})=\coprod_{k\geq 0}\mathcal{S}_{\xi}(k).$ (8)
@@
The objects of $\mathcal{C}_{\xi}$ are pairs $(\sigma,i)$, where $\sigma$ is a contiguous subword of $\xi$, and $i$ is the starting position of this subword in $\xi$ (positions start from 0). For example,
@@
  • $(\texttt{abc},0)$ denotes abc at position 0.
@@
  • $(\texttt{a},0)\to(\texttt{abc},0)$, since a is a prefix of abc starting at 0.
@@
The above diagram shows morphisms between objects of word length 0, 11, and $2$, where objects of length 0 map to those of length $1$, and objects of length $1$ map to those of length $2$. In addition, compositions of morphisms are also morphisms; for example, the morphism $(\texttt{b},3)\to(\texttt{bcb},1)$ is the composition of the morphism $(\texttt{b},3)\to(\texttt{cb},2)$ and the morphism $(\texttt{cb},2)\to(\texttt{bcb},1)$.
@@ 3.2 Distances on resolution category
The Levenshtein distance is a widely used distance that quantifies the minimum number of edit operations (insertions, deletions, and substitutions) required to transform one sequence into another. The LCS distance, as the name suggests, is based on the length of the longest common subsequence shared by two sequences, providing a measure of their structural similarity. Specifically, it is defined as
$d_{\mathrm{LCS}}(X,Y)=1-\frac{\mathrm{LCS}(X,Y)}{\min(\#X,\#Y)},$ (10)
where $\mathrm{LCS}(X,Y)$ denotes the length of the longest common subsequence of sequences $X$ and $Y$. Here, #X\#X denotes the length (i.e., the number of elements) of the sequence $X$.
The Hamming distance, particularly suitable for sequences of equal length, counts the number of positions at which the corresponding elements differ. However, these classical distances primarily focus on the intrinsic differences between two sequences themselves. They are not designed to capture the differences in the occurrences or contextual roles of two sequences within a larger sequence.
@@
$(\sigma_{1},i_{1})\leftarrow(\sigma_{0},i_{0})\rightarrow(\sigma_{2},i_{2})$ (11)
@@
$d_{\mathrm{int}}((\sigma_{1},i_{1}),(\sigma_{2},i_{2})):=1-\frac{L}{\min\{\mathrm{length}(\sigma_{1}),\mathrm{length}(\sigma_{2})\}}.$ (12)
This distance lies in $[0,1]$, and takes the value 0 if and only if the shorter of the two subsequences is fully contained in a common intersection object. It increases as the shared structure between the two objects becomes smaller.
@@
$d_{\mathrm{pos}}((\sigma_{1},i_{1}),(\sigma_{2},i_{2}))=|i_{1}-i_{2}|.$ (13)
@@
$S(\sigma)=\{\,(\sigma,i)\in\mathrm{Ob}(\mathcal{C}_{\xi})\mid\text{$\sigma$ occurs in $\xi$ starting at position $i$}\,\}.$ (14)
@@
$d_{\mathrm{freq}}(\sigma_{1},\sigma_{2})=\left|f_{\xi}(\sigma_{1})-f_{\xi}(\sigma_{2})\right|.$ (15)
@@
$d((\sigma_{1},i_{1}),(\sigma_{2},i_{2}))=\alpha\cdot d_{\text{LCS}}(\sigma_{1},\sigma_{2})+\beta\cdot d_{\text{pos}}((\sigma_{1},i_{1}),(\sigma_{2},i_{2})),$ (16)
@@
Although this section introduces several different distance measures, in our subsequent applications, we primarily adopt the position distance. The position distance better captures the fine-grained features of sequences of moderate length and is well suited for the tasks considered in this work. In comparison, the intersection distance is more appropriate for analyzing relatively short sequences, while the frequency distance is more efficient for sequences with lengths in the tens or even hundreds of millions. The use of the combined distance depends largely on the specific requirements of the task. Naturally, other distance measures can also be considered, as each possesses its own advantages and is suited to different application scenarios.
@@ 3.3 Substructure complexes in the resolution category
Assume that a distance function $d$ is defined on the set of objects of $\mathcal{C}_{\xi}$. Let $U\subseteq\operatorname{Ob}(\mathcal{C}_{\xi})$ be a finite collection of objects. For each real number $r\in\mathbb{R}$, we define a $\Delta$-complex $\mathcal{K}_{r}(U)$ as follows: for $p\geq 1$, a $p$-simplex in $\mathcal{K}_{r}(U)$ is a subset $\{\sigma_{0},\dots,\sigma_{p}\}\subseteq U$ of $p+1$ objects such that $d(\sigma_{i},\sigma_{j})\leq r$ for all $0\leq i,j\leq p$. And the vertex set of $\mathcal{K}_{r}(U)$ is $U/\sim$. Here, $\sim$ represents equivalence of objects at the same position, i.e., $(\sigma_{1},i)\sim(\sigma_{2},i)$ for any $\sigma_{1}$ and $\sigma_{2}$. This complex is called the substructure complex associated with $U$. Substructure complexes of a resolution category are also referred to as resolution complexes.
Consider the case where $U=S(\sigma)$. In this setting, $\mathcal{K}_{r}(U)$ coincides with the Vietoris–Rips complex of the metric subspace $(U,d)$ at scale parameter $r$. We denote by $\mathcal{K}^{n}_{r}(U)$ the set of all $n$-simplices in the complex $\mathcal{K}_{r}(U)$. As $r$ increases, this construction defines a filtration $\{\mathcal{K}_{r}(U)\}_{r\in\mathbb{R}}$, which gives rise to persistent homology. The resulting persistence module captures the multi-scale topological features of the object collection $U$ within $\mathcal{C}_{\xi}$.
@@
$H_{n}^{a,b}(U)\coloneqq{\mathrm{im}\hskip 1.00006pt}\left(f^{a,b}_{n}:H_{n}(\mathcal{K}_{a}(U))\to H_{n}(\mathcal{K}_{b}(U))\right),\quad n\geq 0.$ (18)
@@
$d_{\mathrm{pos}}((\sigma_{1},i_{1}),(\sigma_{2},i_{2}))=|i_{1}-i_{2}|.$ (19)
@@
$\begin{pmatrix}&(C,1)&(G,3)&(C,5)&(G,8)\\ (C,1)&0&2&4&7\\ (G,3)&2&0&2&5\\ (C,5)&4&2&0&3\\ (G,8)&7&5&3&0\\ \end{pmatrix}$ (20)
@@
$H_{n}(\mathcal{K}_{2}(U))=\begin{cases}\mathbb{K}\oplus\mathbb{K},&\text{if }n=0,\\ 0,&\text{otherwise}.\end{cases}$ (21)
@@
$\mathcal{K}^{0}_{4}(U)$ $=\{\{v_{1}\},\{v_{2}\},\{v_{3}\},\{v_{4}\}\},$
$\mathcal{K}^{1}_{4}(U)$ $=\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{1},v_{3}\},\{v_{3},v_{4}\}\},$
$\mathcal{K}^{2}_{4}(U)$ $=\{\{v_{1},v_{2},v_{3}\}\}.$
@@
$H_{n}^{2,4}(U)=\begin{cases}\mathbb{K},&\text{if }n=0,\\ 0,&\text{otherwise}.\end{cases}$ (22)
@@
Let $U=\{A,AG\}$ be the subset of nucleic acid patterns of interest. The positions of $A$ and $AG$ in $\xi$ are as follows: $A$ appears at positions 0, 66, and $7$, and $AG$ appears at position $7$. We associate these occurrences with labeled objects as follows: $(A,0)$, $(A,6)$, $(A,7)$, and $(AG,7)$. For simplicity, we denote these objects as $w_{1}=(A,0)$, $w_{2}=(A,6)$, $w_{3}=(A,7)$, and $w_{4}=(AG,7)$.
@@
$\begin{pmatrix}&(A,0)&(A,6)&(A,7)&(AG,7)\\ (A,0)&0&6&7&7\\ (A,6)&6&0&1&1\\ (A,7)&7&1&0&0\\ (AG,7)&7&1&0&0\\ \end{pmatrix}$ (23)
@@
$\mathcal{K}^{0}_{1}(U)$ $=\{\{w_{1}\},\{w_{2}\},\{w_{3}\}=\{w_{4}\}\},$
$\mathcal{K}^{1}_{1}(U)$ $=\{\{w_{2},w_{3}\},\{w_{2},w_{4}\},\{w_{3},w_{4}\}\},$
$\mathcal{K}^{2}_{1}(U)$ $=\{\{w_{2},w_{3},w_{4}\}\}.$
@@
$H_{n}(\mathcal{K}_{1}(U))=\begin{cases}\mathbb{K}\oplus\mathbb{K},&\text{if }n=0,\\ \mathbb{K},&\text{if }n=1,\\ 0,&\text{otherwise}.\end{cases}$ (24)
@@
$\mathcal{K}^{0}_{6}(U)$ $=\{\{w_{1}\},\{w_{2}\},\{w_{3}\}=\{w_{4}\}\},$
$\mathcal{K}^{1}_{6}(U)$ $=\{\{w_{1},w_{2}\},\{w_{2},w_{3}\},\{w_{2},w_{4}\},\{w_{3},w_{4}\}\},$
$\mathcal{K}^{2}_{6}(U)$ $=\{\{w_{2},w_{3},w_{4}\}\}.$
@@
$H_{n}^{1,6}(U)=\begin{cases}\mathbb{K},&\text{if }n=0,1,\\ 0,&\text{otherwise}.\end{cases}$ (25)
@@
For a given DNA sequence of length $N$, the computational complexity for constructing a $k$-dimensional resolution complex is, in the worst case, $O(kN)$. With the use of optimized algorithms, this complexity can be improved to $O(N)$. For the entire pipeline, the overall computational complexity is typically $O(N\log N)$, dominated by Betti number computations. However, due to the special nature of the position distance, the use of advanced optimization techniques and iterative methods can reduce the complexity to nearly $O(N)$.
@@ 4.1 CTSA for predictions of protein-nucleic acid binding affinities
For nucleic acids, we validate CTSA in DNA sequences to assess its capability to reveal essential interaction patterns relevant to protein binding. CTSA provides a topological realization of DNA by introducing resolution categories that reflect the intricate connections of the sequence within its k-mer spaces. Persistent homology is then applied across these categories to extract topological features, namely, persistent Betti numbers, which are essential for CTSA embeddings to capture both local and global structure of sequence. The resulting CTSA feature vectors have a dimension of 800, which is comparable to the 768-dimensional DNABERT features used in the related SVSBI study.
@@ 5 Concluding remarks
Author Contributions
Jian Liu wrote the first draft, developed the theoretical framework and methods, and revised the manuscript. Shen Li wrote the first draft, designed the algorithms, created the figures, and participated in revision. Mushal Zia contributed to the theoretical analysis, data collection, and algorithm development. Guo-Wei Wei conceived the main ideas, supervised the project, revised the manuscript, and secured funding.
Funding Information
This work was supported in part by NIH grants R01GM126189, R01AI164266, and R35GM148196, National Science Foundation grants DMS2052983, DMS-1761320, and IIS-1900473, NASA grant 80NSSC21M0023, Michigan State University Research Foundation, and Bristol-Myers Squibb 65109. Jian was also supported by Natural Science Foundation of China (NSFC Grant No. 12401080) and Scientific Research Foundation of Chongqing University of Technology.
Conflict of Interest
The authors declare no competing interests.
Statement of Usage of Artificial Intelligence
Artificial intelligence tools or technologies were not used in this project except for limited English grammar corrections.